OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). (sin x)(cos x) = 0 A. pi/2, pi B. 0, pi/2, pi, 3pi/2 C. pi, 3pi/2 D. 0, 3pi/2

OpenStudy (perl):

hint sin x cos x = sin(2x)/2

geerky42 (geerky42):

So either \(\sin x = 0\) or \(\cos x=0\). So you know where in unit circle ( interval [0, 2π) ) is where \(\sin x = 0\) or \(\cos x=0\)?

OpenStudy (perl):

(sin x)(cos x) = 0 has the same solutions as sin(2x)/2 = 0 sin(2x) = 0 x = arcsin(0)/2

OpenStudy (anonymous):

So, sin x = 0?

OpenStudy (perl):

nope

OpenStudy (perl):

sin(2x) = 0 , theres a difference

OpenStudy (anonymous):

Sinx =0 at 0 and also pi and multples of pi. Cosx =0 at pi/2 and odd multiples of pi/2. So the answer is B?